OpenStudy (sleepyhead314):

|x + 2| = -x solve for x

OpenStudy (sleepyhead314):

@iGreen.

OpenStudy (anonymous):

You break it into two equations.. x + 2 = -x -x - 2 = -x

OpenStudy (anonymous):

I g2g now, try and figure it out

OpenStudy (sleepyhead314):

*bonks myself on the head* RIGHT thanks

OpenStudy (mimi_x3):

Lol do you get it ....@iGreen. Kinda simplified it

OpenStudy (sleepyhead314):

I just forgot what absolute values meant for something like that

OpenStudy (mimi_x3):

lol kkk